Two cars pass exit 20 on the turnpike at 8:00 a.m. One car is traveling north and the other is traveling south at a speed that is 4 km/h faster than the northbound car. At 8:45 a.m. the cars are 117 km apart. Find the speed of the southbound car.
Answer by MathTherapy(639)

Two cars pass exit 20 on the turnpike at 8:00 a.m. One car is traveling north and the other is traveling south at a speed that is 4 km/h faster than the northbound car. At 8:45 a.m. the cars are 117 km apart. Find the speed of the southbound car.
.75(N + 4) + .75N = 117
.75N + 3 + .75N = 117
1.5N = 114
N = 76
Since southbound car was going 4 km/h faster than the northbound car, then the southbound car was going (76 + 4) = 80 km/h.